QUESTION 21: The Antenna Reciprocity Theorem states an antenna’s electromagnetic properties will be the same whether it is transmitting or receiving.

True /False

QUESTION 22: Select the correct statement(s) regarding antenna and RF wave polarization.

a. a vertically polarized RF wave cannot be collected by a horizontally polarized antenna

b. right-hand circular polarization is said to be orthogonal to left-hand circular polarization

c. signals of the same frequency can be transmitted together in an non-interfering manner as long as their polarizations are orthogonal to one another

d. all statements are correct

QUESTION 23: Depolarization of a transmitted signal reduces the amount of energy transfer between the RF wave and antenna.

True / False

QUESTION 24: Given the following information, determine the SNR in decibels at the receiver: transmitted power [Pt]=100 dBW, free space loss [FSL]=80 dB, noise [N]= -10 dBW, power received [Pr]=[Pt]-[FSL], [SNR]=[Pr]-[N].

a. [SNR]= 170 dB

b. [SNR]= 10 dB

c. [SNR]= 190 dB

d. [SNR]= 30 dB
Answer #1

21 Answer: true
Explanation: The antenna’s electromagnetic properties like bandwidth, impedance, polarization and resonant frequency are same whether the signals are transmitting or receiving. This refers to the Antenna Reciprocity Theorem.

22 Answer: option d. All statements are correct.

a. The vertically polarised RF wave can be collected by a verticalpolarisedantenna but not collected by a horizontalpolarised antenna.
b. If electric field vector rotates counter anti-clockwise as far as it gets from the antenna, it is called right hand circular polarization.
If electric field vector rotates counter clockwise as far as it gets from the antenna, it is called left hand circular polarization.
Therefore, the right-hand polarization is orthogonal to left hand polarization.
c. The signals of the same frequency can be transmitted together in a non-interfering manner as long as their polarizations are orthogonal to one another.

23 Answer: true
Explanation: Depolarization of a transmitted signal reduces the amount of energy transfer between the RF wave and antenna.

23 Answer: option d

Consider the given data.
The transmitted power [Pt]=100 dBW.
The free space loss [FSL]=80 dB.
The noise [N]= -10 dBW.
The power received [Pr]=[Pt]-[FSL].
[SNR]=[Pr]-[N].

Convert the dbW (decibel Watts) to decibel milliwatt.
The decibel milliwatt is 30 more than the decibel Watts.
Then, transmitted power [Pt]=100 dBW = 130 dBm
The noise [N]= -10 dBW = 20 dBm.
The power received [Pr]=[Pt]-[FSL] = 130 – 80 = 50 db
The value of SNR =[Pr]-[N] = 50 -20 = 30 db.

Therefore, option d is correct.